Operation Summary for October 28, 2011 - October 31, 2011 |
NuMI reported receiving 9.5E18 protons on target for the period from 10/24/11 to 10/31/11. Friday October 28 The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, with the Neutron Therapy Facility (NTF) preparing to treat two patients later in the morning, with Antiproton Source (Pbar) personnel conducting Mu2e and reverse proton studies, and with the NuMI and MiniBooNE experiments taking beam. During the day shift, Operations turned off MI beam at 9:27 AM for one-hour of work in the MI-60 power supply room. Operations established beam to the Meson FTBF experiment T-1015 around 10:40 AM. At 1:49 PM, operators reported a power glitch that tripped off RF stations, power supplies, and beam in Linac, Booster, and MI; the High Voltage Coordinator investigated and reported that it was a ComEd problem with an undetermined cause. Operations reestablished beam at 2 PM. Pbar personnel ended their studies at 2:08 PM. Operators prepared Pbar for access. Operations turned off the Linac beam at 2:57 PM for maintenance work. During the evening shift, at 7:02 PM, the MiniBooNE beam permit tripped off due to beam position problems; it reset. A Linac RF station (LRF1) tripped off for the first of two times in seven minutes; it reset. Operations disabled beam to T-1015 at 10 PM. Saturday October 29 The midnight shift began with Operations sending beam to Switchyard, NuMI, and MiniBooNE. During the midnight shift, operators tuned the Booster and MiniBooNE machines throughout the shift to increase efficiencies and raise beam intensity. During the day shift, at 9:24 AM, a Booster RF station (BRF4) tripped off for the first of multiple times in five minutes; operators adjusted the station’s anode program. Operations established beam to T-1015 at 10:03 AM. At 2:34 PM, a Booster vacuum pump (B:IP241) tripped off for the first of three times in 25 minutes; it reset. During the evening shift, Operations disabled beam to T-1015 at 10:02 PM. At 10:44 PM, operators reported finding a number of vacuum pumps that had tripped off in the 8 GeV transfer line; operators turned off beam to reset the pumps. 8 GeV vacuum had recovered by 11:03 PM. Sunday October 30 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, operators tuned the 750 keV line, 400 MeV line, and MiniBooNE. During the day shift, Operations injected protons into the Recycler at 8:13 AM. Operations established beam to T-1015 at 10:54 AM. At 12:57 PM, Operations received an alarm for Linac; Operations turned off the Linac beam and operators investigated. Operators reported that a Linac Klystron RF station (KRF5) was smoking. Operators called in Linac personnel. Linac repairs ended and Linac resumed sending beam at 3:45 PM. During the evening shift, two I- Source vacuum pumps tripped off at 7:17 PM; everything reset. Operations disabled beam to T-1015 at 10:05 PM. Monday October 31 The midnight shift began with Operations sending beam to NuMI and MiniBooNE. During the midnight shift, operators tuned the 750 KeV line and Booster. Operations Coordinator (OPSCO) The OpsCo approved of the plans by Booster and Recycler personnel to conduct studies. This is the last day that the Antiproton Source will go by that name. Starting tomorrow they will become the Muon Department. |
|
More Information Comments and Suggestions |
| last modified 10/31/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |